# BIV-Priv-Seg: Locating Private Content in Images Taken by People With Visual Impairments

### v2

@@ Abstract
Individuals who are blind or have low vision (BLV) are at a heightened risk of sharing private information if they share photographs they have taken. To facilitate developing technologies that can help preserve privacy, we introduce BIV-Priv-Seg, the first localization dataset originating from people with visual impairments that shows private content. It contains 1,028 images with segmentation annotations for 16 private object categories. We first characterize BIV-Priv-Seg and then evaluate modern models’ performance for locating private content in the dataset. We find modern models struggle most with locating private objects that are not salient, small, and lack text as well as recognizing when private content is absent from an image. We facilitate future extensions by sharing our new dataset with the evaluation server at https://vizwiz.org/tasks-and-datasets/o
@@ 1 Introduction
A concern is that many images from BLV photographers contain private content. Underscoring this concern, one study showed that privacy leaks occurred for over 10% of more than 40,000 images shared by BLV photographers using a visual assistance technology (i.e., the VizWiz application) [18]. Such privacy leaks are concerning regardless of whether they are seen by remote humans providing visual assistance or retained by companies [45, 46], who may subsequently use such data for training AI models and for accruing licensing revenue from other end users.
Figure 1: We introduce BIV-Priv-Seg, a dataset containing segmented private objects. An example is shown from each category.
@@
We also characterize BIV-Priv-Seg and compare it with five existing few-shot localization datasets to reveal what makes it similar to prior work as well as unique. We found that differentiating aspects of BIV-Priv-Seg include no target objects in 19% of the images, objects exhibiting greater variability in terms of how much of an image they occupy, objects residing more often on image borders, and objects more frequently containing text.
@@ 2 Related Work
A critical component for successfully developing computer vision models is the availability of human-annotated datasets to support reproducible, quantitative evaluation. A challenge for developing privacy-aware technologies is that most datasets lack private content due to intentional filtering efforts, including for popular, mainstream datasets such as VOC [12], ImageNet/ILSVRC [41], COCO [29], LSUN [51], and Open Images [27]. Only relatively small collections of private images (e.g., thousands of examples) have been successfully curated for public use, including by scraping them from photo-sharing websites (i.e., Flickr, Twitter) [36, 37, 53] and hiring BLV photographers to photograph (non-personal) private content [44]. Extending the latter prior work, we introduce the first publicly-available dataset originating from BLV photographers that shows segmented private content. This effort extends the broader research community’s inclusivity efforts to increase the availability of datasets based on images taken by people with vision impairments [3, 6, 8, 9, 19, 20, 25, 31, 17, 18, 52].
Given that it is non-trivial to curate a large-scale collection of private images for supporting the development of privacy-aware models, since the very nature of such data means it typically cannot be saved or shared, we focus on introducing a dataset challenge that facilitates the development of zero-shot and few-shot models (and so requires a small amount of training data). Accordingly, our dataset challenge will complement prior work that proposed few-shot object detection and instance segmentation challenges, including VizWiz-FewShot [47], PASCAL-$5^{i}$ [43], COCO-$20^{i}$ [35, 32], and LVIS [34], and FSOD [14]. Our dataset offers unique challenges compared to existing datasets, including images lacking the target object, a wider variability of object sizes, a higher prevalence of objects residing on image borders, and a higher prevalence of objects with text.
@@
Zero-shot and few-shot methods that can locate objects in images span more traditional few-shot object localization models and more recent vision-language models. Traditional methods include those designed to only achieve strong performance on novel object categories [38, 32, 10, 50, 28, 30, 21, 11, 15, 33, 34] as well as those designed to also maintain performance on base object categories [4, 16] (i.e., generalized/incremental setting). A new breed of relevant models also emerged in the past year: large multimodal models (LMMs) that can generate a textual description alongside a visual grounding (i.e., segmentation mask) specifying where the described content is located [39, 49]. [^1] Also related is a framework introduced in 2024 [42] for adapting vision-language models (VLMs) to identify privacy levels of conversational scenarios. However, their work does not support locating private content, restricting its scope for potential privacy-preserving technologies. LMMs offer the advantage of being more interactive, making them particularly appealing for assistive tools for the BLV community. We benchmark a total of three models (2 traditional and 1 LMM) to assess how well modern methods perform on our new dataset. While we observe the best performance from LMMs, all models still fall below human performance. Our results suggest that promising future directions for research are to improve models’ abilities to handle objects that are not salient, small, and lack text as well as to recognize when private content is absent from an image.
@@ 3.1 Dataset Creation
We leverage the BIV-Priv dataset [44], which consists of 1,176 images taken by 26 BLV photographers of 16 types of private objects. All private objects were ‘props’, meaning that the objects lacked any private information pertinent to the photographer (e.g., a fake medical bill rather than a real one). Each BLV individual took pictures with each object type positioned to appear in the image’s foreground as well as its background. We leverage all images, including those that lacked the target private content, to reflect the unique challenges faced by BLV photographers who often do not successfully capture the content of interest in their images
@@
We chose $16$ private object categories, after slightly modifying the original categories proposed in BIV-Priv [44] (as discussed in the Supplementary Materials). Our final set of categories is as follows: “bank statement”, “bill or receipt”, “business card”, “condom box”, “condom packet”, “credit or debit card”, “doctor’s prescription”, “pill bottle”, “letter with address”, “local newspaper”, “medical record document”, “mortgage or investment report”, “pregnancy test”, “pregnancy test box”, “tattoo sleeve”, and “transcript”.
@@
We leveraged the instructions and annotation protocol established by prior work (i.e., VizWiz-FewShot [47]) to enable fair comparison with it. To create instance segmentations, annotators clicked a series of points to generate polygons, dragged a polygon’s vertices to refine it, and created holes by drawing ‘interior’ polygons in ‘exterior’ polygons. We leveraged the annotation platform, Supervise.Ly [^2] https://supervisely.com/ to collect all segmentation annotations.
@@
We next localized all instances of the 16 private object categories in every image. To collect high-quality annotations, we relied on three in-house annotators who were authors of this work and so intimately familiar with the images. Each individual had a 1-hour 1-on-1 training session to learn about the annotation interface and then passed a test demonstrating they could accurately annotate a sample image with the annotation tool. Each image was annotated by two annotators and we used $IoU$ scores to determine how to establish a ground truth segmentation per image. When $IoU\geq 0.8$, we randomly chose one annotation for the ground truth. Otherwise, the three authors reviewed the pair of annotations to choose one as the ground truth (or, in exceptional cases, discarded both annotations). [^3] Only $15\%$ of the images required manual review, with $93\%$ having one of the annotations deemed correct. This result culminated in 967 instance segmentations.
Figure 2: Examples from our BIV-Priv-Seg dataset illustrating its unique aspects, specifically a high variability for the proportion of images occupied by the objects (with many objects occupying much of their images’ borders), high prevalence of text in objects, and inclusion of images lacking target objects.
@@ 3.2 Dataset Analysis
We analyze three properties for each few-shot localization dataset: the number of annotated objects per image, the number of disconnected areas an object contains (i.e., segments), and the percentage of objects containing text [^4] Objects are flagged if Microsoft Azure’s Optical Character Recognition (OCR) API detects text within the image, after masking out all content except for the instance segmentation.. PASCAL-5i and FSOD are excluded from the latter two metrics since their annotations only contain bounding boxes and so lack the necessary segmentation information to count segments and locate texts. Results are visualized in Figure 3.
@@
When observing the histogram showing the percentage of images containing different numbers of annotated objects (Figure 3(a)), we observe two key distinctions of our dataset. First, our dataset is the only one with images lacking objects of interest, with this occurring for 18% of the images. These reflect both when no target object is visible and when the target object is a paper positioned on its blank side, rather than the side showing the document’s text (and so private text as well). This amount is similar to that reported by prior work [9], where roughly 20% of BLV photographers’ images lacked the visual evidence needed to answer their visual question despite showing recognizable content [20]. We retain such images to reflect that the photographers intended to capture an object of interest in such images [44], highlighting the challenge of dealing with target-less samples. This is to encourage modern algorithms to achieve a good true negative rate, which is critical for many real-world applications. To support fair dataset comparisons, we exclude these images when analyzing our dataset for the rest of this section. The second key distinction of our dataset is that it rarely has more than one annotated object per image. We attribute this to the distinct photography goal for BIV-Priv [44], where the goal for every image was to photograph a single object, whereas other datasets such as COCO-20i were intentionally curated to show complex scenes showing more than one object (i.e., none-iconic images).
@@
When observing the histogram showing the percentage of images containing different numbers segments (i.e., disconnected areas) (Figure 3(b)), we observe that our BIV-Priv-Seg dataset typically exhibits only one segment per object. VizWiz-FewShot, in contrast, more often has the most with up to $122$ segments in an object. While an object itself exists as a unified entity in the real world, multiple counts of segments arise in an object under two circumstances: (1) when occlusions break objects into more than one region and (2) when objects are partially outside the frame of the image so there are disconnected regions of an object. Our dataset limits these scenarios, thereby permitting more focused model benchmarking analysis for other challenges that can arise, as discussed in this section.
@@
When analyzing the prevalence of text in objects, we find it is the highest for our dataset, occurring for $74.5\%$ of annotated object instances (Figure 3(c)). We attribute this greater prevalence of text to the intrinsic nature of the private categories, where textual information often is the private content. For instance, paper-based categories such as “bill or receipt” and “doctor’s prescriptions” can show names, addresses, and more. Consequently, our dataset provides an opportunity to shift algorithm developers’ focus to more heavily prioritize analyzing text on objects.
@@
We next analyze the properties of annotated object instances. We apply metrics to each dataset’s annotations as is, whether they are segmentations or bounding boxes. For every instance, we compute the following four metrics:
@@
    Boundary complexity: one minus the isoperimetric inequality, which is the ratio of the area of an instance to the length of its perimeter. A lower isoperimetric inequality value indicates a more complex boundary. This metric is informative only for segmentation annotations, so we exclude datasets with only bounding box annotations.
@@ 4.1 Few-shot Localization Algorithms
We benchmarked the top-performing few-shot object detection (FSOD) and few-shot instance segmentation (FSIS) models for which publicly available code could be successfully deployed on modern GPUs: DeFRCN [38] for FSOD and YOLACT [7] for both FSOD and FSIS. [^5] We adopted the hyperparameters and training pipelines from the prior work [38] for DeFRCN and the prior work [47] for YOLACT. We trained the models for base categories using the larger-scale VizWiz-FewShot [47], since those examples also originate from BLV individuals, and then fine-tuned the models on the categories in BIV-Priv-Seg.
@@
The experiments were conducted using 4-fold cross-validation, where the 100 categories in VizWiz-FewShot were divided into four sets, with the combination of three sets (75 categories) serving as base categories in each fold. This setting matches prior work [47] for a fair comparison to VizWiz-FewShot.
@@
We evaluated the models using the mean Average Precision (mAP), which is the average precision across a range of intersection over union (IoU) thresholds from 0.5 to 0.95 in steps of 0.05. This metric aligns with prior work [47, 38, 33, 22]. We also present results with respect to $mAP_{50}$, where only threshold $0.5$ is used since this approach facilitates the comparison with datasets such as Pascal VOC.
@@
For comparison, we also analyzed each trained model’s performance for other few-shot localization datasets: PASCAL-$5^{i}$[43], COCO-20i[35, 32], and VizWiz-FewShot [47].
Overall Performance.
Results are in Tables 1666The benchmarking results reported for DeFRCN performance on PASCAL-$5^{i}$ is presented in the format specified in [43] with three data splits. To compare these results to the 4-fold cross-validation format, we average the scores of the three pre-defined splits. and 2.
Overall, we observe similar DeFRCN performance on our dataset to that observed for the baseline datasets (Table 1). With respect to the mAP scores, our dataset appears to be easier compared to VizWiz-FewShot and COCO-20i, while it is harder than PASCAL-5i when considering the AP50 scores. A commonality across all the datasets is that the DeFRCN’s performance still is quite poor, falling short of what might be practically useful in an end-user-facing tool. Overall, these findings underscore the potential benefit of our dataset in helping advance modern few-shot learning algorithms.
@@
FSOD mAP FSOD AP50 FSIS mAP FSIS AP50
Dataset   \   $k=$ 1 3 5 10 1 3 5 10 1 3 5 10 1 3 5 10
COCO-20i [35] - - - - 19.0 - 21.7 - - - - - 17.1 - 18.9 -
Figure 5: Qualitative results from DeFRCN in the 1-shot setting for object detection (bounding boxes), YOLACT in the 1-shot setting for instance segmentation (masks) and object detection (bounding boxes), and GLaMM using the prompt I scenario for the zero-shot setting. As a point of reference, ground truth results are also provided.
Qualitative results from the 1-shot settings are shown for both models in Figure 5. In these examples, we can see that YOLACT and DeFRCN tend to miss the same objects, while YOLACT is able to locate some objects that DeFRCN misses.
@@
We next perform fine-grained analysis of the models by dividing the test examples into subsets based on the following characteristics also employed by prior work [47]: object size and presence of text. We categorize objects into small, medium, and large sizes, by sorting all instances in BIV-Priv-Seg based on their image coverage and then equally divide them into three groups. To detect the presence of text, we used OCR on background-masked instance images to flag whether text is present. Results are provided in Table 3 with respect to mAP scores. The trends of AP50 scores are consistent with that of the mAP scores and so are excluded from the table.
For object size, our findings for YOLACT reinforce those of prior work [23, 54, 47], showing that it performs worse for smaller objects. However, DeFRCN’s results indicate that the models tend to perform better on medium-sized objects. The observation of these differences in model trends provides an insight for future BLV applications that the unique object size characteristics could play a crucial role in the design of these models.
With respect to presence of text, we observed that both models consistently perform better on objects with text than without text. This finding supports the observations in previous work [47] that texts could serve as strong features for the models to localize objects in few-shot learning. This emphasizes the advantages of localization of private objects based on the high prevalence of text, highlighting a strong foundation for privacy applications in the area.

### v3

@@ Abstract
Individuals who are blind or have low vision (BLV) are at a heightened risk of sharing private information if they share photographs they have taken. To facilitate developing technologies that can help them preserve privacy, we introduce BIV-Priv-Seg, the first localization dataset originating from people with visual impairments that shows private content. It contains 1,028 images with segmentation annotations for 16 private object categories. We first characterize BIV-Priv-Seg and then evaluate modern models’ performance for locating private content in the dataset. We find modern models struggle most with locating private objects that are not salient, small, and lack text as well as recognizing when private content is absent from an image. We facilitate future extensions by sharing our new dataset with the evaluation server at https://vizwiz.org/tasks-and-datasets/o
@@ 1 Introduction
A concern is that many images from BLV photographers contain private content [1]. Underscoring this concern, one study showed that privacy leaks occurred for over 10% of more than 40,000 images shared by BLV photographers using a visual assistance technology (i.e., the VizWiz application) [20]. Such privacy leaks are concerning regardless of whether they are seen by remote humans providing visual assistance or retained by companies [49, 50], who may subsequently use such data for training AI models and for accruing licensing revenue from other end users.
Figure 1: We introduce BIV-Priv-Seg, a dataset containing segmented private objects. An example is shown for each category.
@@
We also characterize BIV-Priv-Seg and compare it with five existing few-shot localization datasets to reveal what makes it similar to prior work as well as unique. Similarities to prior findings about images from blind photographers include that objects exhibit greater variability in terms of how much of an image they occupy and objects reside more often on image borders. Differentiating aspects of BIV-Priv-Seg include the lack of presence of target objects in 19% of the images, a high rate of images containing one annotated object, and a higher prevalence of objects containing text.
@@ 2 Related Work
A critical component for successfully developing computer vision models is the availability of human-annotated datasets to support reproducible, quantitative evaluation. A challenge for developing privacy-aware technologies is that most datasets lack private content due to intentional filtering efforts, including popular, mainstream datasets such as VOC [13], ImageNet/ILSVRC [46], COCO [32], LSUN [56], and Open Images [30]. Only relatively small collections of private images (e.g., thousands of examples) have been successfully curated for public use, including by scraping them from photo-sharing websites (i.e., Flickr, Twitter) [40, 41, 58] and hiring BLV photographers to photograph (non-personal) private content [48]. Extending the latter prior work, we introduce the first publicly-available dataset originating from BLV photographers that shows segmented private content. This effort extends the broader research community’s inclusivity efforts to increase the availability of datasets based on images taken by people with vision impairments [3, 6, 9, 10, 21, 22, 27, 35, 19, 20, 57] while following the standard practice of augmenting annotations to images that are already used in other datasets [52, 18, 15, 32].
Given that it is non-trivial to curate a large-scale collection of private images for supporting the development of privacy-aware models, since the very nature of such data means it typically cannot be saved or shared, we focus on introducing a dataset challenge that facilitates the development of zero-shot and few-shot models (and so requires a small amount of training data). Accordingly, our dataset challenge will complement prior work that proposed few-shot object detection and instance segmentation challenges, including VizWiz-FewShot [52], PASCAL-$5^{i}$ [47], COCO-$20^{i}$ [39, 36], LVIS [38], and FSOD [15]. Our dataset offers unique challenges compared to existing datasets, including images lacking the target object, a wider variability of object sizes, a higher prevalence of objects residing on image borders, and a higher prevalence of objects with text.
@@
Zero-shot and few-shot methods that can locate objects in images span more traditional few-shot object localization models and more recent vision-language models. Traditional methods include those designed to only achieve strong performance on novel object categories [42, 36, 11, 55, 31, 34, 23, 12, 16, 37, 38] as well as those designed to also maintain performance on base object categories [4, 17] (i.e., generalized/incremental setting). A new breed of relevant models also emerged in the past year: large multimodal models (LMMs) that can generate a textual description alongside a visual grounding (i.e., segmentation mask) specifying where the described content is located [44, 54]. LMMs offer the advantage of being more interactive, making them particularly appealing for assistive tools for the BLV community. We benchmark a total of three models (2 traditional and 1 LMM) to assess how well modern methods perform on our new dataset. While we observe the best performance from LMMs, all models still fall below human performance. Our results suggest that promising future directions for research are to improve models’ abilities to handle objects that are not salient, small, and lack text as well as to recognize when private content is absent from an image.
@@ 3.1 Dataset Creation
We leverage the BIV-Priv dataset [48], which consists of 1,176 images taken by 26 BLV photographers of 16 types of private objects. All private objects were ‘props’, meaning that the objects lacked any private information pertinent to the photographer (e.g., a fake medical bill rather than a real one). Each BLV individual took pictures with each object type positioned to appear in the image’s foreground as well as its background. We leverage all images, including those that lacked the target private content, to reflect the unique challenges faced by BLV photographers who often do not successfully capture the content of interest in their images.
@@
After slightly modifying the categories in BIV-Priv [48], as discussed in the Supplementary Materials, we chose $16$ private object categories as follows: “bank statement", “bill or receipt", “business card", “condom box", “condom packet", “credit or debit card", “doctor’s prescription", “pill bottle", “letter with address", “local newspaper", “medical record document", “mortgage or investment report", “pregnancy test", “pregnancy test box", “tattoo sleeve", and “transcript".
@@
We leveraged the instructions and annotation protocol established by prior work (i.e., VizWiz-FewShot [52]) to enable fair comparison with it. To create instance segmentations, annotators clicked a series of points to generate polygons, dragged a polygon’s vertices to refine it, and created holes by drawing ‘interior’ polygons in ‘exterior’ polygons. We leveraged the annotation platform, Supervise.Ly (i.e., https://supervisely.com/) to collect all segmentation annotations.
@@
We next localized all instances of the 16 private object categories in every image. To collect high-quality annotations, we relied on three in-house annotators who were authors of this work and so intimately familiar with the images (matching the approach employed for creating ADE20K [60]). Each individual had a 1-hour 1-on-1 training session to learn about the annotation interface and then passed a test demonstrating they could accurately annotate a sample image. Each image was annotated by two annotators and we used $IoU$ scores to determine how to establish a ground truth segmentation per image (matching the approach employed for COCO-Stuff [8]). When $IoU\geq 0.8$, we randomly chose one annotation for the ground truth. Otherwise, the three authors reviewed the pair of annotations to choose one as the ground truth (or, in exceptional cases, discarded both annotations). Only $15\%$ of the images required manual review, with $93\%$ having one of the annotations deemed correct. This effort culminated in 967 instance segmentations.
Figure 2: Examples from BIV-Priv-Seg dataset illustrating its unique aspects: (a) high proportion of image areas and borders occupied by the objects, (b) high prevalence of objects containing text, and (c) images lacking target objects.
@@ 3.2 Dataset Analysis
We analyze three properties for each few-shot localization dataset: the number of annotated objects per image, the number of disconnected areas an object contains (i.e., segments), and the percentage of objects containing text [^1] Objects are flagged as containing text if Microsoft Azure’s Optical Character Recognition (OCR) API detects text within the image, after masking out all content except for the instance segmentation.. PASCAL-5i and FSOD are excluded from the latter two metrics since their annotations only contain bounding boxes and so lack the necessary segmentation information to count segments and locate texts. Results are visualized in Figure 3.
@@
From the histogram showing the percentage of images containing each number of annotated objects (Figure 3(a)), we observe two key distinctions of our dataset. First, our dataset is the only one with images lacking objects of interest with it occurring for 18% of the images. These reflect both when no target object is visible and when the target object is a paper positioned on its blank side, rather than the side showing the document’s text (and so the private text). This amount is similar to that reported by prior work [10], where roughly 20% of BLV photographers’ images lacked the visual evidence needed to answer their visual question despite showing recognizable content [22]. We retain such images in the final dataset to reflect that the photographers intended to capture an object of interest in such images [48], but exclude these examples for subsequent data analysis to support fair dataset comparisons. Retaining these target-less samples will facilitate developing models that recognize when no target objects are present, which is relevant for many real-world applications.
The second key distinction of our dataset based on Figure 3(a) is that it rarely has more than one annotated object per image. This trend aligns with findings from prior work [20], which show that only one private object is typically observed in an image when privacy leaks arise for BLV photographers authentically trying to learn about their visual surroundings. We attribute this finding in our dataset to the distinct photography goal for BIV-Priv [48], where images were intentionally photographed with a single private object. However, we suspect many images in BIV-Priv show multiple objects, because when evaluating the performance of a salient object detection model [28] (shown to yield comparable performance to humans for images taken by people with visual impairments [45]), we found that only 32% of the annotated private objects overlap significantly (IoU $\geq 0.5$) with salient areas. The presence of multiple objects could be in part due to how images were curated [48], as photographers were instructed to position private objects in the background for half the images.
@@
When observing the histogram showing the percentage of images containing different numbers of segments (i.e., disconnected areas) (Figure 3(b)), we observe that our BIV-Priv-Seg dataset typically exhibits only one segment per object. VizWiz-FewShot, in contrast, more often has the most with up to $122$ segments in an object. While an object itself exists as a unified entity in the real world, multiple counts of segments arise in an object under two circumstances: (1) when occlusions break objects into more than one region and (2) when objects are partially outside the frame of the image so there are disconnected regions of an object. Our dataset limits these scenarios, thereby permitting more focused model benchmarking analysis for other challenges that can arise, as discussed in this section.
@@
When analyzing the prevalence of text in objects, we find it is the highest for our dataset, occurring for $74.5\%$ of annotated object instances (Figure 3(c)). We attribute this greater prevalence of text to the intrinsic nature of the private categories, where textual information often is the private content. For instance, paper-based categories such as “bill or receipt" and “doctor’s prescriptions" can show names, addresses, and more. Consequently, our dataset provides an opportunity to shift algorithm developers’ focus to more heavily prioritize analyzing text on objects.
FSOD mAP FSOD AP50 FSIS mAP FSIS AP50
Dataset   \   $k=$ 1 3 5 10 1 3 5 10 1 3 5 10 1 3 5 10
COCO-20i [39] - - - - 19.0 - 21.7 - - - - - 17.1 - 18.9 -
@@
We next analyze the annotated object instances, characterizing each dataset’s annotations as is for both segmentations and bounding boxes. For every instance, we compute the following:
@@
    Boundary complexity: one minus the isoperimetric inequality, which is the ratio of the area of an instance to the length of its perimeter. A lower isoperimetric inequality value indicates a more complex boundary. This metric cannot be computed for bounding box annotations, so we exclud such datasets.
@@ 4.1 Few-shot Localization Algorithms
We benchmarked the top-performing few-shot object detection (FSOD) and few-shot instance segmentation (FSIS) models for which publicly available code could be successfully deployed on modern GPUs: DeFRCN [42] for FSOD and YOLACT [7] for both FSOD and FSIS. We adopted the hyperparameters and training pipelines from [42] for DeFRCN and [52] for YOLACT. We trained models for base categories on the larger-scale VizWiz-FewShot [52], since those examples also originate from BLV individuals, and then fine-tuned them for novel categories on BIV-Priv-Seg.
@@
The experiments were conducted using 4-fold cross-validation, where the 100 categories in VizWiz-FewShot were divided into four sets, with the combination of three sets (75 categories) serving as base categories in each fold. This matches prior work [52] to enable fair comparison with VizWiz-FewShot [52].
@@
We evaluated the models using mean Average Precision (mAP), which is the average precision across a range of intersection over union (IoU) thresholds from 0.5 to 0.95 in steps of 0.05. This metric aligns with prior work [52, 42, 37, 24]. We also present results with $AP_{50}$, where only threshold $0.5$ is used to enable comparison with benchmarking on datasets such as Pascal VOC.
@@
We also analyzed each model’s performance on other few-shot datasets: PASCAL-$5^{i}$[47], COCO-20i[39, 36], and VizWiz-FewShot [52].
@@
Figure 5: Qualitative results from DeFRCN in the 1-shot setting for object detection, YOLACT in the 1-shot setting for instance segmentation and object detection, and GLaMM and GroundingDINO+SAM using the prompt I scenario for the zero-shot setting.
Overall Performance.
Results are in Tables 1 and $2$The benchmarking results reported for DeFRCN performance on PASCAL-$5^{i}$ is presented in the format specified in [47] with three data splits. To compare these results to the 4-fold cross-validation format, we average the scores of the three pre-defined splits..
@@
Overall, we observe similar DeFRCN performance on our dataset to that observed for the baseline datasets (Table 2). With respect to the mAP scores, our dataset appears to be easier compared to VizWiz-FewShot and COCO-20i, while it is harder than PASCAL-5i when considering the AP50 scores. A commonality across all the datasets is that the DeFRCN’s performance is still quite poor, falling short of what might be practically useful in an end-user-facing tool. Overall, our findings underscore the potential benefit of our dataset in helping advance modern few-shot learning algorithms.
Qualitative results from the 1-shot settings are shown for both models in Figure 5. In these examples, we can see that YOLACT and DeFRCN tend to miss the same objects, while YOLACT locates some objects that DeFRCN misses.
@@
We next perform fine-grained analysis of the models by dividing the test examples into subsets based on the following characteristics also employed by prior work [52]: object size and presence of text. We categorize objects into small, medium, and large sizes, by sorting all instances in BIV-Priv-Seg based on their image coverage and then equally divide them into three groups. Results are provided in Table 3 with respect to mAP scores. The trends of AP50 scores are consistent with that of the mAP scores and so are excluded from the table.
For object size, both model tend to perform better for larger objects, which is promising given that this size is a distinctly prominent feature in images from BLV photographers [45, 9]. However, DeFRCN overall performs best on medium-sized objects rather than large-sized objects, suggesting a possibly slightly inferior match for BLV photographers. Both models perform the worst for smaller objects, reinforcing findings from prior work [25, 59, 52].
With respect to presence of text, we observed that both models consistently perform better when objects have text. This reinforces findings in previous work [52], that text could serve as a valuable feature for models to locate objects in few-shot learning. This is particularly important for our privacy detection setting, since there is such a strong correlation between private content and the presence of text.